OpenStudy (anonymous):

what is the equation for a line perpendicular to the line 3x+2y=6 passing through (1,-2)

OpenStudy (anonymous):

A perpedicular line to L has a gradient equal to the negative reciprocal of that of L. L:3x+2y-6=0 => N:2x-3y+c=0 Plugging in (1,-2) as a point on N we get 2+6+c=0 +>c=-8 N:2x-3y=8

OpenStudy (anonymous):

Write 3x + 2y = 6 in standard form........... y = mx + c 2y = -3x + 6 then devide both sides by 2 y= -3/2x + 3.........................(i) mi = -3/2 but m of perpendicular line is - reciprocal of mi, thus 2/3 so y = 2/3x + c to get c we substitute values of 1,-2) in the equation -2 = 2/3(1) + c c = -2 - 2/3............... =-5/3 substituting -5/3 with,we have y = 2/3x +(-5/3) y = 2/3x - 5/3 as the equation of perpendicular line.